### v1

@@
Abstract
Accurate, high-resolution, and real-time DOA estimation is a cornerstone of environmental perception in automotive radar systems. While sparse signal recovery techniques offer super-resolution and high-precision estimation, their prohibitive computational complexity remains a primary bottleneck for practical deployment. This paper proposes a sparse DOA estimation scheme specifically tailored for the stringent requirements of automotive radar such as limited computational resources, restricted array apertures, and single snapshot. By introducing the concept of the spatial angular pseudo-derivative and incorporating this property as a constraint into a standard $\ell_{0}$-norm minimization problem, we formulate an objective function that more faithfully characterizes the physical properties of the DOA problem. The associated solver, designated as the SAPD search algorithm, naturally transforms the high-dimensional optimization task into an efficient grid-search scheme. The SAPD algorithm circumvents high-order matrix inversions and computationally intensive iterations. We provide an analysis of the computational complexity and convergence properties of the proposed algorithm. Extensive numerical simulations demonstrate that the SAPD method achieves a superior balance of real-time efficiency, high precision, and super-resolution, making it highly suitable for next-generation automotive radar applications.
{IEEEkeywords}
DOA estimation, Sparse Signal Recovery, Spatial Angular Pseudo-Derivative, Single Snapshot, Computation Efficiency.
1 INTRODUCTION
@@
While super-resolution direction of arrival (DOA) estimation has been extensively studied for over four decades [15], [16], achieving real-time accuracy remains a formidable challenge under stringent constraints such as limited computational resources, restricted array apertures, and single snapshot. Notably, this technology is pivotal for automotive millimeter-wave (mmWave) radars, where it enables critical perception tasks within autonomous driving systems and robot environment [29], [13], [31], [10], [7]. Therefore, addressing these challenges to enhance the angular estimation performance of mmWave radars is essential to meet the requirements [28] of next-generation autonomous driving.
@@
In autonomous driving scenarios, the DOA estimation process must be executed for all range-Doppler (RD) cells [4] identified by Constant False Alarm Rate (CFAR) detection within the strict duration of a single frame interval. Typically, the frame interval of automotive mmWave radars is less than 50 ms, during which multiple RD cells must be processed [17], [36]. Considering the time overhead incurred by post-processing following point cloud generation, computational efficiency becomes the primary prerequisite for the practicality of any super-resolution DOA algorithm. Failure to achieve effective DOA estimation within the required time frame inevitably results in the loss of critical target information, thereby compromising the integrity of the generated point clouds and severely undermining the reliability of the downstream radar processing pipeline. In addition, estimation accuracy and angular resolution are equally vital metrics for evaluating super-resolution DOA algorithms. Consequently, a super-resolution DOA algorithm with practical potential needs simultaneously provide low computational complexity, high accuracy, and superior angular resolution [28].
@@
Taking radar simultaneous localization and mapping (SLAM) as a representative post-processing scenario, DOA estimation typically relies on spatial discrete Fourier transforms (DFT) or conventional beamforming (CBF) techniques, such as Bartlett beamforming, to satisfy real-time constraints [10], [7], [12]. CBF is capable of performing DOA estimation with only a single snapshot and exhibits inherent robustness to coherent signals. However, the angular resolution of these methods is fundamentally constrained by the array aperture, a physical boundary known as the Rayleigh limit [22]. In many DOA estimation studies [15], [16], [38], [5], [8], when the angular separation of two sources falls within the Rayleigh limit, the CBF spatial spectrum is generally considered to have failed in resolution due to peak merging. Consequently, the information embedded within such spectra is rarely further exploited. Recently, Itzhak et al. [11] proposed a maximum array gain beamforming method designed to minimize signal distortion across a predefined region-of-interest (ROI). While this approach more effectively leverages information within the spatial spectrum by accounting for all directions within the ROI, it is primarily intended for signal enhancement and does not provide explicit DOA estimates. In addition to limited angular resolution, the estimation accuracy of CBF is inferior to that of super-resolution algorithms. This lack of precision exacerbates the accumulation of errors within the SLAM framework, thereby complicating downstream processing tasks [7], [19].
@@
To bridge the gap between sparse signal recovery theory and practical implementation, this paper proposes a sparse DOA estimation scheme tailored for meet the real-time, high-accuracy, and high-resolution demands of automotive radar systems. The crux of the proposed method lies in the discovery that the off-grid model, in conjunction with the spatial grid, exhibits a property analogous to a first-order derivative relative to the true DOAs. This characteristic is termed the spatial angular pseudo-derivative. By leveraging this property, we reformulate the sparse DOA estimation problem, transforming it from a computationally intensive high-dimensional optimization into an efficient grid search procedure. We define and characterize the spatial angular pseudo-derivative, a mathematical property that endows discrete grid points with directional information relative to the true targets. The main contributions of this work are summarized as follows:
@@
  • We develop the computationally efficient SAPD Search algorithm. By leveraging the properties of the spatial angular pseudo-derivative, this algorithm circumvents high-order matrix inversions and intensive iterative procedures. The process begins with initialization using the spatial spectrum obtained via Bartlett beamforming, followed by a targeted search for the ground-truth DOAs. Finally, a greedy patch mechanism is integrated to address complex scenarios and recover potential missed detections in the initial support set, thereby ensuring real-time, high-precision, and high-resolution DOA estimation for automotive radar.
  • An analysis of the computational complexity and convergence properties of the SAPD algorithm are provided.
  • Extensive numerical experiments validate that the proposed method achieves superior performance in terms of real-time operation, high precision, and high resolution for automotive radar DOA.
@@
Notations: $\|\cdot\|_{2}$, $\|\cdot\|_{1}$ denote the $\ell_{2}$-norm and $\ell_{1}$-norm, respectively. $(\cdot)^{T}$, $(\cdot)^{H}$, $(\cdot)^{-1}$ and $(\cdot)^{\dagger}$ represent the transposition, Hermitian transposition, inversion, pseudo-inversion, respectively. $\lceil\cdot\rceil$ and $\operatorname{round}(\cdot)$ denote the ceiling and rounding operators, respectively. $\mathbb{E}(\cdot)$ denotes the expectation operator. $j=\sqrt{1}$. $\mathbf{1}_{N}$ is the $N\times 1$ all-ones vector. $n]=\{1,\dots n\}$.
2 Problem Statement and Preliminary Analysis
@@
where $\mathbf{R}_{\boldsymbol{y}}$ denotes the signal covariance matrix. In the single-snapshot case, this expression simplifies to $P_{\text{g}}(\theta)=|\boldsymbol{a}^{H}(\theta)\boldsymbol{y}|^{2}$. We define the vectorized spatial spectrum as $P_{\text{g}}(\boldsymbol{\theta})\in\mathbb{R}^{G}$, where each element corresponds to the power estimated at the respective grid point. The spatial spectrum encapsulates significant information regarding the spatial distribution of the targets, providing the basis for the initialization of our proposed scheme.
@@
$\boldsymbol{x}_{\mathcal{G}}$ $=[x_{g_{1}},x_{g_{2}},\dots,x_{g_{d}}]^{T},\ g_{i}\in\mathcal{G},$ (7)
@@
3 Spatial Angular Pseudo-Derivative Search Algorithm
@@
3.1 Proposed Objective and spatial angular pseudo-derivative concept
@@
$\boldsymbol{\beta}(\boldsymbol{\theta}^{*}-\Delta\boldsymbol{\theta})>0,\ \boldsymbol{\beta}(\boldsymbol{\theta}^{*}+\Delta\boldsymbol{\theta})<0,$ (14)
@@
$(P_{D})\begin{aligned} &\ \mathop{\arg\min}\limits_{\mathcal{G}}\|\boldsymbol{y}-\mathbf{A}(\boldsymbol{\theta}_{\mathcal{G}})\boldsymbol{x}(\boldsymbol{\theta}_{\mathcal{G}})\|_{2}\\ \ \ \text{s.t.}\ |\mathcal{G}|=&K,\ \boldsymbol{\beta}(\boldsymbol{\theta}_{\mathcal{G}}-\Delta\boldsymbol{\theta})>0,\ \boldsymbol{\beta}(\boldsymbol{\theta}_{\mathcal{G}}+\Delta\boldsymbol{\theta})<0.\end{aligned}$ (15)
@@
$h(\boldsymbol{\theta}_{\mathcal{G}})=$ $\operatorname{sng}{(\boldsymbol{\beta}(\boldsymbol{\theta}_{\mathcal{G}}-\Delta\boldsymbol{\theta}))}+\operatorname{sng}{(\boldsymbol{\beta}(\boldsymbol{\theta}_{\mathcal{G}}+\Delta\boldsymbol{\theta}))}.$ (17)
@@
$\text{s.t.}\quad|\mathcal{G}|=K,$ $\ h(\boldsymbol{\theta}_{\mathcal{G}})=0.$
@@ 3.2 SAPD Scheme
3.3 Computational complexity analysis
@@
3.4 Convergence Analysis of SAPD Search
@@
4 Simulation Results
@@
In Fig.5LABEL:sub@Throughput_CT, we evaluate the computational throughput by varying the number of detected RD-cells from 1 to 50. The algorithms are executed sequentially, and the total processing latency for successful DOA estimation is recorded. Although OMP exhibits shorter execution times, it is excluded from this comparison due to its insufficient estimation precision, which fails to provide meaningful super-resolution results. The RMSE of OMP represents in the Table 1.
@@
As illustrated in Fig.5LABEL:sub@Throughput_CT, for a single RD-cell, the proposed SAPD algorithm is the only method capable of providing valid estimates in under 1 ms. While IAA-APES and VALSE also satisfy the 50-ms frame interval constraint, the remaining benchmarks exceed this threshold. When the workload increases to 10 RD-cells, SAPD remains the sole algorithm operating within the real-time budget, requiring less than 10 ms. For the most demanding case of 50 RD-cells, SAPD completes the estimation in 36.4 ms, effectively maintaining real-time feasibility within the 50-ms window, whereas all other compared methods require 1000 ms or more. The RMSE performance of the algorithms capable of completing a single RD-cell estimation within 50 ms, namely OMP, IAA-APES, VALSE, and the proposed SAPD, is summarized in Table 1. Among these compared algorithms, the proposed SAPD search scheme achieves the minimum RMSE.
Table 1: The RMSE of four algorithm in the scenario
@@
RMSE $4.8988^{\circ}$ $0.2130^{\circ}$ $0.3453^{\circ}$ $\boldsymbol{0.2035^{\circ}{}^{*}}$
@@
In view of real-time processing constraints, Table 2 restricts the RMSE comparison to computationally efficient methods, specifically VALSE, IAA-APES, and the proposed SAPD. Under the specified simulation settings, VALSE fails to resolve the targets, yielding invalid estimates. While IAA-APES achieves an RMSE of $0.8944^{\circ}$, the proposed SAPD algorithm demonstrates superior precision with a significantly lower RMSE of $0.2724^{\circ}$, substantially outperforming the IAA-APES benchmark.
@@
Table 2: The RMSE of three algorithm in the scenario
@@
RMSE - $0.8944^{\circ}$ $\boldsymbol{0.2724^{\circ}{}^{*}}$
Table 3 summarizes the execution times for the three algorithms. Notably, VALSE incurs a 140ms latency even in cases of estimation failure, whereas IAA-APES requires 30ms. In stark contrast, the proposed SAPD algorithm achieves a processing time of merely 1.242 ms, inclusive of the Greedy Patch mechanism overhead.
Table 3: The computational time of three algorithm in the scenario
@@
Computation Time $140\mathrm{ms}$ $30\mathrm{ms}$ $\boldsymbol{1.242\mathrm{ms}{}^{*}}$
@@
5 Conclusion
This paper introduced the spatial angular pseudo-derivative framework, a sparse DOA estimation scheme specifically designed for single-snapshot, high-precision, and real-time applications under stringent computational constraints. By leveraging the concept of the spatial angular pseudo-derivative, we transformed the high-dimensional sparse recovery task into an efficient search-based solver. Both theoretical analysis and numerical results indicate that the proposed scheme achieves sub-millisecond execution times without compromising estimation fidelity. Notably, its resolution capability exceeds that of several high-complexity methods like DML. The practical effectiveness of the SAPD algorithm was further verified using raw radar data collected from a hardware platform. Given its superior performance and minimal computational footprint, the SAPD algorithm represents a promising solution for the sensing requirements of next-generation autonomous driving.
[…]
6 Proof of Theorem 1

### v2

@@
Spatial Angular Pseudo-Derivative Searching: A Single Snapshot Super-resolution Sparse DOA Scheme with Potential for Practical Application
Longxin Bai, Jingchao Zhang, and Liyan Qiao This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible. This work is supported by National Natural Science Foundation of China under Grant 61701138, the Natural Science Foundation of Heilong jiang Province of China under Grant LH2022F019, Hei Long Jiang Postdoctoral Foundation under Grant LBH-Z16087 and the Fundamental Research Funds for the Central Universities under Grant HIT.NSRIF202339. (Corresponding author: Jingchao Zhang) L. Bai, J. Zhang, and L. Qiao are with School of Electronics and Information Engineering, Harbin Institute of Technology, Harbin 150001, China.(e-mail: bailongxin@stu.hit.edu.cn, zhangjingchao@hit.edu.cn, qiaoliyan@hit.edu.cn)
I INTRODUCTION
@@
While super-resolution direction-of-arrival (DOA) estimation has been extensively studied for over four decades [15], [16], achieving real-time accuracy remains a formidable challenge under stringent constraints such as limited computational resources, restricted array apertures, and the single-snapshot constraint. Notably, this technology is pivotal for automotive millimeter-wave (mmWave) radars, where it enables critical perception tasks within autonomous driving systems and robotic environments [29], [13], [31], [10], [7]. Therefore, addressing these challenges to enhance the angular estimation performance of mmWave radars is essential to meet the requirements [28] of next-generation autonomous driving.
@@
In autonomous driving scenarios, the DOA estimation process must be executed for all Range-Doppler (RD) cells [4] identified by Constant False Alarm Rate (CFAR) detection within the strict duration of a single frame interval. Typically, the frame interval of automotive mmWave radars is less than 50 ms, during which multiple RD cells must be processed [17], [36]. Considering the time overhead incurred by post-processing following point cloud generation, computational efficiency becomes the primary prerequisite for the practicality of any super-resolution DOA algorithm. Failure to achieve effective DOA estimation within the required time frame inevitably results in the loss of critical target information, thereby compromising the integrity of the generated point clouds and severely undermining the reliability of the downstream radar processing pipeline. In addition, estimation accuracy and angular resolution are equally vital metrics for evaluating super-resolution DOA algorithms. Consequently, a super-resolution DOA algorithm with practical potential needs to simultaneously provide low computational complexity, high accuracy, and superior angular resolution [28].
I-A Prior work
@@
Taking radar simultaneous localization and mapping (SLAM) as a representative post-processing scenario, DOA estimation typically relies on spatial Discrete Fourier Transforms (DFT) or conventional beamforming (CBF) techniques, such as Bartlett beamforming, to satisfy real-time constraints [10], [7], [12]. CBF is capable of performing DOA estimation with only a single snapshot and exhibits inherent robustness to coherent signals. However, the angular resolution of these methods is fundamentally constrained by the array aperture, a physical boundary known as the Rayleigh limit [22]. In many DOA estimation studies [15], [16], [38], [5], [8], when the angular separation of two sources falls within the Rayleigh limit, the CBF spatial spectrum is generally considered to fail to resolve the sources due to peak merging. Consequently, the information embedded within such spectra is rarely further exploited. Recently, Itzhak et al. [11] proposed a maximum array gain beamforming method designed to minimize signal distortion across a predefined region of interest (ROI). While this approach more effectively leverages information within the spatial spectrum by accounting for all directions within the ROI, it is primarily intended for signal enhancement and does not provide explicit DOA estimates. In addition to limited angular resolution, the estimation accuracy of CBF is inferior to that of super-resolution algorithms. This lack of precision exacerbates the accumulation of errors within the SLAM framework, thereby complicating downstream processing tasks [7], [19].
@@
I-B Contribution
To bridge the gap between sparse signal recovery theory and practical implementation, this paper proposes a sparse DOA estimation scheme tailored to meet the real-time, high-accuracy, and high-resolution requirements of automotive radar systems. The crux of the proposed method lies in the discovery that the off-grid model, in conjunction with the spatial grid, exhibits a property analogous to a first-order derivative relative to the true DOAs. This characteristic is termed the spatial angular pseudo-derivative. By leveraging this property, we reformulate the sparse DOA estimation problem, transforming it from a computationally intensive high-dimensional optimization into an efficient grid search procedure. We define and characterize the spatial angular pseudo-derivative, a mathematical property that endows discrete grid points with directional information relative to the true targets. The main contributions of this work are summarized as follows:
@@ I-B Contribution
  • We propose the computationally efficient SAPD search algorithm. By leveraging the properties of the spatial angular pseudo-derivative, this algorithm circumvents large-scale matrix inversions and intensive iterative procedures. The process begins with initialization using the spatial spectrum obtained via Bartlett beamforming, followed by a targeted search for the ground-truth DOAs. A greedy patch mechanism is integrated to address complex scenarios and recover potential missed detections in the initial support set, thereby ensuring real-time, high-precision, and high-resolution DOA estimation for automotive radar systems.
  • An analysis of the computational complexity and convergence properties of the SAPD algorithm is provided.
  • Numerical simulations and experimental validation demonstrate that the proposed method achieves superior performance in terms of real-time operation, high precision, and high resolution for automotive radar DOA estimation.
@@
Notations: $\|\cdot\|_{2}$, $\|\cdot\|_{1}$ denote the $\ell_{2}$-norm and $\ell_{1}$-norm, respectively. $(\cdot)^{T}$, $(\cdot)^{H}$, $(\cdot)^{-1}$ and $(\cdot)^{\dagger}$ represent the transposition, Hermitian transposition, inversion, pseudo-inversion, respectively. $\lceil\cdot\rceil$ and $\operatorname{round}(\cdot)$ denote the ceiling and rounding operators, respectively. $\mathbb{E}(\cdot)$ denotes the expectation operator. $j=\sqrt{1}$. $\mathbf{1}_{N}$ is the $N\times 1$ all-ones vector. $n]=\{1,\dots n\}$. $\operatorname{sgn}(\cdot)$ denotes the signum function.
II Problem Statement and Preliminary Analysis
@@
where $\mathbf{R}_{\boldsymbol{y}}$ denotes the observation data covariance matrix. In the single-snapshot case, this expression simplifies to $P_{\text{g}}(\theta)=|\boldsymbol{a}^{H}(\theta)\boldsymbol{y}|^{2}$. We define the vectorized spatial spectrum as $P_{\text{g}}(\boldsymbol{\theta})\in\mathbb{R}^{G}$, where each element corresponds to the power estimated at the respective grid point. The spatial spectrum encapsulates significant information regarding the spatial distribution of the targets, providing the basis for the initialization of our proposed scheme.
@@
$\boldsymbol{x}_{\mathcal{G}}$ $=[x_{g_{1}},x_{g_{2}},\dots,x_{g_{d}}]^{T},\ g_{i}\in\mathcal{G}.$ (7)
@@
III Spatial Angular Pseudo-Derivative Search Algorithm
@@
III-A Proposed Objective and spatial angular pseudo-derivative concept
@@
$\boldsymbol{\beta}(\boldsymbol{\theta}^{*}-\Delta\boldsymbol{\theta})>\mathbf{0},\ \boldsymbol{\beta}(\boldsymbol{\theta}^{*}+\Delta\boldsymbol{\theta})<\mathbf{0},$ (14)
@@
$(P_{D})\begin{aligned} &\ \mathop{\arg\min}\limits_{\mathcal{G}}\|\boldsymbol{y}-\mathbf{A}(\boldsymbol{\theta}_{\mathcal{G}})\boldsymbol{x}(\boldsymbol{\theta}_{\mathcal{G}})\|_{2}\\ \ \ \text{s.t.}\ |\mathcal{G}|=&K,\ \boldsymbol{\beta}(\boldsymbol{\theta}_{\mathcal{G}}-\Delta\boldsymbol{\theta})>\mathbf{0},\ \boldsymbol{\beta}(\boldsymbol{\theta}_{\mathcal{G}}+\Delta\boldsymbol{\theta})<\mathbf{0}.\end{aligned}$ (15)
@@
$h(\boldsymbol{\theta}_{\mathcal{G}})=$ $\operatorname{sgn}{(\boldsymbol{\beta}(\boldsymbol{\theta}_{\mathcal{G}}-\Delta\boldsymbol{\theta}))}+\operatorname{sgn}{(\boldsymbol{\beta}(\boldsymbol{\theta}_{\mathcal{G}}+\Delta\boldsymbol{\theta}))}.$ (17)
@@
$\text{s.t.}\quad|\mathcal{G}|=K,$ $\ h(\boldsymbol{\theta}_{\mathcal{G}})=\mathbf{0}.$
@@ III-B SAPD Scheme
III-C Computational complexity analysis
@@
III-D Convergence Analysis of SAPD Search
@@
IV Numerical Simulations and Experimental Validation
@@
In Fig.5LABEL:sub@Throughput_CT, we evaluate the computational throughput by varying the number of detected RD-cells from 1 to 50. The algorithms are executed sequentially, and the total processing latency for successful DOA estimation is recorded. Although OMP exhibits shorter execution times, it is excluded from this comparison due to its insufficient estimation precision, which fails to provide meaningful super-resolution results. The RMSE of OMP represents in the Table I.
@@
As illustrated in Fig.5LABEL:sub@Throughput_CT, for a single RD-cell, the proposed SAPD algorithm is the only method capable of providing valid estimates in under 1 ms. While IAA-APES and VALSE also satisfy the 50-ms frame interval constraint, the remaining benchmarks exceed this threshold. When the workload increases to 10 RD-cells, SAPD remains the sole algorithm operating within the real-time budget, requiring less than 10 ms. For the most demanding case of 50 RD-cells, SAPD completes the estimation in 36.4 ms, effectively maintaining real-time feasibility within the 50-ms window, whereas all other compared methods require 1000 ms or more. The RMSE performance of the algorithms capable of completing a single RD-cell estimation within 50 ms, namely OMP, IAA-APES, VALSE, and the proposed SAPD, is summarized in Table I. Among these compared algorithms, the proposed SAPD search scheme achieves the minimum RMSE.
TABLE I: The RMSE of four algorithm in the scenario
@@
RMSE $4.8988^{\circ}$ $0.2130^{\circ}$ $0.3453^{\circ}$ $\boldsymbol{0.2035^{\circ}}$
@@
In view of real-time processing constraints, Table II restricts the RMSE comparison to computationally efficient methods, specifically VALSE, IAA-APES, and the proposed SAPD. Under the specified simulation settings, VALSE fails to resolve the targets, yielding invalid estimates. While IAA-APES achieves an RMSE of $0.8944^{\circ}$, the proposed SAPD algorithm demonstrates superior precision with a significantly lower RMSE of $0.2724^{\circ}$, substantially outperforming the IAA-APES benchmark.
@@
TABLE II: The RMSE of three algorithm in the scenario
@@
RMSE - $0.8944^{\circ}$ $\boldsymbol{0.2724^{\circ}}$
Table III summarizes the execution times for the three algorithms. Notably, VALSE incurs a 140ms latency even in cases of estimation failure, whereas IAA-APES requires 30ms. In stark contrast, the proposed SAPD algorithm achieves a processing time of merely 1.242 ms, inclusive of the Greedy Patch mechanism overhead.
((a))
((b))
((a))
((b))
TABLE III: The computational time of three algorithm in the scenario
@@
Computation Time $140\mathrm{ms}$ $30\mathrm{ms}$ $\boldsymbol{1.242\mathrm{ms}}$
@@
V Conclusion
This paper introduced the spatial angular pseudo-derivative framework, a sparse DOA estimation scheme specifically designed for single-snapshot, high-precision, and real-time applications under stringent computational constraints. By leveraging the concept of the spatial angular pseudo-derivative, we transformed the high-dimensional sparse recovery task into an efficient search-based solver. Theoretical analysis demonstrates that the proposed algorithm has low computational complexity, and numerical results indicate that the proposed scheme achieves millisecond execution times without compromising estimation fidelity. Notably, its resolution capability exceeds that of several high-complexity methods like DML. The practical effectiveness of the SAPD algorithm was further verified using raw radar data collected from a hardware platform. Given its superior performance and minimal computational footprint, the SAPD algorithm represents a promising solution for the sensing requirements of next-generation autonomous driving.
Appendix A Proof of Theorem 1